OpenStudy (anonymous):

How do you solve rational equations? 1/2+1/4=1/x I need someone to explain it to me.

OpenStudy (anonymous):

add up on the left \[\frac{1}{2}+\frac{1}{4}=\frac{3}{4}\] start with \[\frac{3}{4}=\frac{1}{x}\] then flip both sides to solve for \(x\)

OpenStudy (mathstudent55):

Another method is to multiply both sides by the LCD. The LCD of 2, 4, and x is 4x: \[(4x)(\frac{ 1 }{ 2 } + \frac{ 1 }{ 4 }) = (4x)(\frac{ 1 }{ x })\] \[(4x)(\frac{ 1 }{ 2 }) + (4x)(\frac{ 1 }{ 4 }) = (4x)(\frac{ 1 }{ x })\] 2x + x = 4 Now continue from here.